OpenStudy (anonymous):

I really need help with this question.... Darian's IRA has fallen 12.6% since the beginning of the year to 142,500.49. What was the value of her IRA at the beginning of the year?

OpenStudy (anonymous):

143500.49 x .126 =18081.06 18081.06+143500.49=$161581.55 is your answer

OpenStudy (anonymous):

Thank you so much

OpenStudy (dumbcow):

Assume your initial amount(I) is 100% if you decrease by 12.6% 100-12.6 = 87.4% Then 142,500.49 = 0.874*I Just divide by .874 to find I

OpenStudy (dumbcow):

same answer but fewer steps :)

OpenStudy (anonymous):

what do you mean by finding 1

OpenStudy (dumbcow):

oh sorry its just the variable i picked to represent the initial amount at beginning of year. call it whatever you want

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Okay I have another question... Translate the following situation into an equation. do not solve. On monday afternoon kelly withdrew 5% of the balance of her banking account to loan her friend some money. On tuesday morning she deposited $100 into her account. She now has $25 more than she did on monday morning. How much was in her account Monday morning?